# Support Graph Preconditioners for Off-Lattice Cell-Based Models

### v1

@@ Abstract
Off-lattice agent-based models (or cell-based models) of multicellular systems are increasingly used to create in-silico models of in-vitro and in-vivo experimental setups of cells and tissues, such as cancer spheroids, neural crest cell migration, and liver lobules. These applications, which simulate thousands to millions of cells, require robust and efficient numerical methods. At their core, these models necessitate the solution of a large friction-dominated equation of motion, resulting in a sparse, symmetric, and positive definite matrix equation. The conjugate gradient method is employed to solve this problem, but this requires a good preconditioner for optimal performance.
In this study, we develop a graph-based preconditioning strategy that can be easily implemented in such agent-based models. Our approach centers on extending support graph preconditioners to block-structured matrices. We prove asymptotic bounds on the condition number of these preconditioned friction matrices. We then benchmark the conjugate gradient method with our support graph preconditioners and compare its performance to other common preconditioning strategies.
1 Introduction
@@
Agent-based models that simulate individual entities such as humans, animals, or biological cells are an indispensable tool for studying emergent behaviors in complex systems. Over the last few decades, biomedical research has adopted agent-based models to develop digital-twins of in-vitro and in-vivo experiments on cell cultures and tissues [5]. To capture a wide variety of applications and research questions, many different agent-based models have been developed. Broadly, we categorize them into lattice-based (e.g. Cellular Automata [18] or Cellular Potts models [15]) and off-lattice models [25]. These different models each have advantages and disadvantages. For an overview, we refer the reader to the review by [38]. Here, we focus on off-lattice models closely related to colloidal physics [10]. In these models, cells are approximated by elastic spheroids [11, 14], capsules [6], ellipsoids [31], or surfaces of triangulated meshes [25]. The applications of these models are highly varied, including slime-mold aggregation [31], cancer growth and migration [23, 27], cancer monolayers and spheroids [11, 39], liver lobules [20], and neural crest cells [29].
@@
Figure 1: Graphical overview of our preconditioner construction. The figure illustrates the step-by-step construction of our proposed preconditioner, proceeding from left to right. 1. Agent configuration. The initial setup showing individual agents (cells) in a spatial arrangement. 2. The collision graph. Using collision detection algorithms, we construct a graph where nodes represent agents and edges represent collisions or friction interactions between them. The friction matrix is the graph laplacian of the collision graph. 3. The maximum spanning tree. Using Prim’s algorithm, we construct a maximum spanning tree from the collision graph. The graph laplacian of this tree, is used as the preconditioner.
@@
A typical cell configuration of spherical cells with radii $R_{i}$ is shown in Fig. 1. Advancing the simulation from $t\to t+\Delta t$ requires solving the overdamped equation of motion $\Gamma\mathbf{v}=\mathbf{F}$, where $\Gamma$ is the friction matrix, $\mathbf{v}$ the cells’ velocities, and $\mathbf{F}$ the forces. The matrix $\Gamma$ is block-structured, symmetric, and positive definite because the individual $3\times 3$ friction matrices are symmetric and positive definite [25]. There is a nonzero off-diagonal block in $\Gamma$ at position $(i,j)$ when cells $i$ and $j$ are in contact. We ultimately solve this large linear system using the conjugate gradient method.
@@
Profiling our simulation software used in [25, 39, 6] shows that solving the equation of motion is often the most time-intensive step. We hypothesize that this is because no good preconditioners are available. From an implementation point of view, it is convenient to implement the linear algebra methods in a matrix-free manner. This means that “off the shelf” preconditioning techniques are difficult to use or adapt. Further, the condition number of the friction matrix $\Gamma$ is determined by the cells’ free surface area. This means that the condition number of the friction matrix varies during a given simulation. Additionally, the sparsity structure of $\Gamma$ is dynamic because it encodes interacting cell pairs. This makes selecting a preconditioner difficult. Here, we solve this by using the collision graph constructed during the collision detection phase as our central data structure instead of the usual sparse matrix implementations.
@@
1.1 Support Graph Preconditioners
The computational cost per iteration of the conjugate gradient method is dominated by the matrix-vector product. Thus, we aim to reduce the number of required iterations. Let $\mathbf{e}_{k}$ denote the true error i.e. the difference between the computed approximation at the $k$-th iteration and the true solution. The well-known error estimate for the conjugate gradient method is given by [34]:
$\norm{\mathbf{e}_{k}}_{\Gamma}\leq 2\left(\frac{\sqrt{\kappa(\Gamma)}-1}{\sqrt{\kappa(\Gamma)}+1}\right)^{k}\norm{\mathbf{e}_{0}}_{\Gamma},$
@@ 1.1 Support Graph Preconditioners
The error estimate suggests that a matrix with higher condition number requires more iterations. However, in practice, conjugate gradient convergence is more complex and often faster than this estimate suggests [26]. While this estimate has limited practical use, it does motivate the reduction of the condition number through preconditioning.
To precondition the system, we choose a symmetric matrix $H=EE^{T}$ such that $\Gamma^{-1}\sim H$. We then solve the modified system $E^{T}\Gamma E\hat{\mathbf{x}}=E^{T}\mathbf{F}$ and $\mathbf{x}=E\hat{\mathbf{x}}$. This approach reduces the iteration count if $\kappa(E^{T}\Gamma E)<\kappa(\Gamma)$, and reduces computational time if we choose $H$ so that solving $H\mathbf{x}=\mathbf{b}$ is computationally inexpensive.
@@
In off-lattice agent-based models, we can interpret the off-diagonal sparsity pattern of $\Gamma$ as a graph $\mathfrak{C}$ (see Fig. 1). This graph representation works as follows:
@@
With employ a technique pioneered by Vaidya [37] that uses subgraphs of $\mathfrak{C}$ as preconditioners. Subgraphs are advantageous because they are sparse, yet capture much of the relevant information from $\mathfrak{C}$. This characteristic allows them to effectively balance between approximating $\Gamma$ and computational efficiency. In our implementation, we specifically use a maximum spanning tree. This tree can be constructed in linearithmic time, and its associated matrix can be factored in linear time. The broader study of using subgraphs as preconditioners is known as support graph theory.
Vaidya’s original manuscript lacked many proofs, which were later provided by [3]. In our work, we extend these proofs to apply to block-structured matrices. This extension allows us to obtain estimates for the smallest and largest eigenvalues of such preconditioned systems. These eigenvalue bounds serve two important purposes: (1) They provide worst-case convergence estimates; and (2) They are valuable in implementing the robust conjugate gradient stopping criteria, proposed by [1, 30].
1.2 Outline
The remainder of this paper is organized as follows: In Section 2, we introduce the linear system arising from agent-based models and its natural graph structure. This section provides the foundation for understanding the mathematical framework of our approach. Section 3 focuses on Vaidya’s preconditioners. We explain how to construct these preconditioners for block-structured matrices and solve the resulting systems in near-linear time. This section bridges the gap between graph theory and numerical linear algebra. Section 4 presents our main theoretical contribution. Here, we extend support graph theory to block-structured matrices and derive eigenvalue bounds for the preconditioned linear system $\Gamma$. This extension is crucial for applying support graph theory to the matrices arising in agent-based models. In Section 5, we present our numerical results. We demonstrate the effectiveness of our preconditioner using a series of numerical benchmarks. These experiments validate our theoretical findings and showcase the practical benefits of our approach. Finally, Section 6 concludes the paper with a discussion of our findings, and its implications for real-world use cases. We situate our contributions in the broader theoretical landscape and discuss connections to the existing literature on this problem. Additionally, we provide potential directions for future research in this area.
2 Preliminaries
@@ 1. Step 1:
    Broad-phase collision detection. Identify possible cell contact pairs $(i,j)$. Efficient divide-and-conquer algorithms such as axis-aligned bounding boxes, are commonly used [36].
@@ 3. Step 3:
    Assemble the friction matrices. Construct the cell-cell and cell-substrate friction matrices. The cell-cell friction matrix between cells $i$ and $j$ is given by
    $\Gamma_{ij}^{cc}=A_{ij}\left(\gamma_{\parallel}\mathbf{u}_{ij}\otimes\mathbf{u}_{ij}+\gamma_{\perp}\left(I-\mathbf{u}_{ij}\otimes\mathbf{u}_{ij}\right)\right),$ (2.1)
    where $\gamma_{\parallel}$ and $\gamma_{\perp}$ are the parallel and perpendicular coefficients of friction, respectively (both of which are positive), and $\mathbf{u}_{ij}$ is the unit contact vector [25]. If $\mathbf{r}_{i}$ is the position of cell $i$, then
    $\mathbf{u}_{ij}=\frac{\mathbf{r}_{j}-\mathbf{r}_{i}}{\norm{\mathbf{r}_{j}-\mathbf{r}_{i}}}.$
@@ 4. Step 4:
    $\Gamma^{cs}_{i}\mathbf{v}_{i}+\sum_{A_{ij}>0}\Gamma^{cc}_{ij}(\mathbf{v}_{i}-\mathbf{v}_{j})=\mathbf{F}_{i},$ (2.2)
    where $\mathbf{v}_{i}$ is the velocity vector and $\mathbf{F}_{i}$ is the total nonfrictional force acting on the cell. We interpret Equation (2.2) as one row of a large linear system $\Gamma\mathbf{v}=\mathbf{F}$. We show that the friction matrix $\Gamma$ is symmetric positive definite. Hence, the conjugate gradient method is an efficient choice for obtaining an accurate solution.
@@ 5. Step 5:
    where the step-size $\Delta t$ is chosen according to the Euler’s method stability criterion. Higher-order integration methods are rarely used. Two exceptions are PhysiCell [14], which employs a second-order Adam’s-Bashford method with a fixed time-step, and [6], where an embedded Runga-Kutta-23 method with an adaptive time-step is employed.
Since the friction matrix $\Gamma$ is comprised of $3\times 3$ blocks, we establish a few simple properties of these smaller friction matrices. Note that each cell-cell friction matrix (and the cell-substrate friction matrix of an ellipsoidal cell) is the sum of two orthogonal projectors. Let $\gamma_{\mathrm{max}}$ and $\gamma_{\mathrm{min}}$ be the maximum and minimum elements of $\{\gamma_{\parallel},\gamma_{\perp}\}$, respectively.
Lemma 2.1
@@
  4. 4.
    its operator norm (with respect to the 2-norm) is $\norm{\Upsilon}=\gamma_{\mathrm{max}}$; and
  5. 5.
@@
Proof: The projection matrices $\mathbf{u}\otimes\mathbf{u}$ and $I-\mathbf{u}\otimes\mathbf{u}$ are clearly symmetric. Note that $(\mathbf{u}\otimes\mathbf{u})\mathbf{u}=\mathbf{u}$, so $\Upsilon\mathbf{u}=\gamma_{\parallel}\mathbf{u}$. Since the eigenvectors of a real symmetric matrix are orthogonal, take $\mathbf{w}$ such that $\mathbf{w}^{T}\mathbf{u}=0$. Then $(\mathbf{u}\otimes\mathbf{u})\mathbf{w}=0$, so $\Upsilon\mathbf{w}=\gamma_{\perp}\mathbf{w}$. The vector $\mathbf{w}$ lies in the orthogonal complement of $\mathbf{u}$, which is two-dimensional. Both of the eigenvalues are positive, which gives positive definiteness. The operator norm of a symmetric positive definite matrix is the maximum eigenvalue. $\blacksquare$
@@
Definition 2.2
A matrix-weighted graph $G=(V,E,w)$ is an undirected graph with a matrix-valued weight function $w\colon V\times V\to\mathbb{R}^{d\times d}$. We require that $w(e)$ is positive definite for all $e\in E$, that $w(v,v)$ is positive semidefinite for all $v\in V$, and that $w(u,v)=0$ otherwise.
@@
Definition 2.3
A block Laplacian is a symmetric block matrix whose off-diagonal blocks are either zero or negative definite, and whose row and column sums are positive semidefinite. For a matrix-weighted graph $G=(V,E,w)$, the block Laplacian of $G$ is the matrix $L$ where
@@
Proof: Let $A$ be a block Laplacian. For all $\mathbf{x}$,
@@
This implies that each term in the summation above is nonnegative, which yields the desired result. $\blacksquare$
3 Support Graph Preconditioners

### v2

@@ Abstract.
Off-lattice agent-based models (or cell-based models) of multicellular systems are increasingly used to create in-silico models of in-vitro and in-vivo experimental setups of cells and tissues, such as cancer spheroids, neural crest cell migration, and liver lobules. These applications, which simulate thousands to millions of cells, require robust and efficient numerical methods. At their core, these models necessitate the solution of a large friction-dominated equation of motion, resulting in a sparse, symmetric, and positive definite matrix equation. The conjugate gradient method is employed to solve this problem, but this requires a good preconditioner for optimal performance. In this study, we develop a graph-based preconditioning strategy that can be easily implemented in such agent-based models. Our approach centers on extending support graph preconditioners to block-structured matrices. We prove asymptotic bounds on the condition number of these preconditioned friction matrices. We then benchmark the conjugate gradient method with our support graph preconditioners and compare its performance to other common preconditioning strategies.
Key words and phrases:
Support graph preconditioner, agent-based model, maximum spanning tree, block Laplacian, conjugate gradient method
2020 Mathematics Subject Classification:
05C50, 65F08, 92-08
1. Introduction
@@
Agent-based models that simulate individual entities such as humans, animals, or biological cells are an indispensable tool for studying emergent behaviors in complex systems. Over the last few decades, biomedical research has adopted agent-based models to develop digital-twins of in-vitro and in-vivo experiments on cell cultures and tissues [4]. To capture a wide variety of applications and research questions, many different agent-based models have been developed. Broadly, we categorize them into lattice-based (e.g. Cellular Automata [18] or Cellular Potts models [15]) and off-lattice models [26]. These different models each have advantages and disadvantages. For an overview, we refer the reader to the review [42]. Here, we focus on off-lattice models closely related to colloidal physics [9]. In these models, cells are approximated by elastic spheroids [10, 14], capsules [5], ellipsoids [33], or surfaces of triangulated meshes [26]. The applications of these models are highly varied, including slime-mold aggregation [33], cancer growth and migration [23, 28], cancer monolayers and spheroids [10, 41], liver lobules [20], and neural crest cells [30].
@@
Figure 1. Graphical overview of our preconditioner construction. The figure illustrates the step-by-step construction of our proposed preconditioner, proceeding from left to right. A) Agent configuration. The initial setup showing individual agents (cells) in a spatial arrangement. B) The collision graph. Using collision detection algorithms, we construct a graph where nodes represent agents and edges represent collisions or friction interactions between them. The friction matrix is the graph Laplacian of the collision graph. C) The maximum spanning tree. Using Prim’s algorithm, we construct a maximum spanning tree from the collision graph. The graph Laplacian of this tree, is used as the preconditioner.
@@
A typical cell configuration of spherical cells with radii $R_{i}$ is shown in Figure 1. Advancing the simulation from $t\to t+\Delta t$ requires solving the overdamped equation of motion $\Gamma\mathbf{v}=\mathbf{F}$, where $\Gamma$ is the friction matrix, $\mathbf{v}$ the cells’ velocities, and $\mathbf{F}$ the forces. The matrix $\Gamma$ is block-structured, symmetric, and positive definite because the individual $3\times 3$ friction matrices are symmetric and positive definite [26]. There is a nonzero off-diagonal block in $\Gamma$ at position $(i,j)$ when cells $i$ and $j$ are in contact. We ultimately solve this large linear system using the conjugate gradient method.
@@
Profiling our simulation software used in [5, 26, 41] shows that solving the equation of motion is often the most time-intensive step. We hypothesize that this is because no good preconditioners have been identified or developed so far. From an implementation point of view, it is convenient to implement the linear algebra methods in a matrix-free manner. This means that “off the shelf” preconditioning techniques are difficult to use or adapt. Further, the condition number of the friction matrix $\Gamma$ is determined by the cells’ free surface area. This means that the condition number of the friction matrix varies during a given simulation. Additionally, the sparsity structure of $\Gamma$ is dynamic because it encodes interacting cell pairs. This makes selecting a preconditioner difficult. Here, we solve this by using the collision graph constructed during the collision detection phase as our central data structure instead of the usual sparse matrix implementations.
@@
1.1. Support Graph Preconditioners
The computational cost per iteration of the conjugate gradient method is dominated by the matrix-vector product. Thus, we aim to reduce the number of required iterations. Let $\mathbf{e}_{k}$ denote the true error at the $k$-th iteration (i.e. the difference between the computed approximation at the $k$-th iteration and the true solution). The well-known error estimate for the conjugate gradient method is given by [36]:
$\lVert\mathbf{e}_{k}\rVert_{\Gamma}\leq 2\left(\frac{\sqrt{\kappa(\Gamma)}-1}{\sqrt{\kappa(\Gamma)}+1}\right)^{k}\lVert\mathbf{e}_{0}\rVert_{\Gamma},$
@@ 1.1. Support Graph Preconditioners
The error estimate suggests that a matrix with higher condition number requires more iterations. However, conjugate gradient convergence is more complex and often faster than this estimate suggests [27]. While this estimate has limited practical use, it does motivate the reduction of the condition number through preconditioning.
To precondition the system, we choose a symmetric matrix $H=EE^{T}$ such that $\Gamma^{-1}\sim H$. We then solve the modified system $E^{T}\Gamma E\hat{\mathbf{x}}=E^{T}\mathbf{F}$ and $\mathbf{x}=E\hat{\mathbf{x}}$. This approach often reduces the iteration count if $\kappa(E^{T}\Gamma E)<\kappa(\Gamma)$, and reduces run time if we choose $H$ so that solving $H\mathbf{x}=\mathbf{b}$ is computationally inexpensive.
@@
In off-lattice agent-based models, we can interpret the off-diagonal sparsity pattern of $\Gamma$ as a graph $\mathfrak{C}$ (see Figure 1). This graph representation works as follows:
@@
We employ a technique pioneered by Vaidya [40] that uses subgraphs of $\mathfrak{C}$ as preconditioners. Subgraphs are advantageous because they are sparse, yet capture much of the relevant information from $\mathfrak{C}$. This characteristic allows them to effectively balance between approximating $\Gamma$ and computational efficiency. In our implementation, we specifically use a maximum spanning tree. This tree can be constructed in linearithmic time, and its associated matrix can be factored in linear time. The broader study of using subgraphs as preconditioners is known as support graph theory.
Vaidya’s original manuscript lacked many proofs, which were later provided in [2]. In our work, we extend these proofs to apply to block-structured matrices. This extension allows us to obtain estimates for the smallest and largest eigenvalues of such preconditioned systems. These eigenvalue bounds serve two important purposes: (1) they provide worst-case convergence estimates, and (2) They are valuable in implementing the robust conjugate gradient stopping criteria proposed by [1, 31].
1.2. Outline
The remainder of this paper is organized as follows: In Section 2, we introduce the linear system arising from agent-based models and its natural graph structure. This section provides the foundation for understanding the mathematical framework of our approach. Section 3 focuses on Vaidya’s preconditioners. We explain how to construct these preconditioners for block-structured matrices and solve the resulting systems in near-linear time. This section bridges the gap between graph theory and numerical linear algebra. Section 4 presents our main theoretical contribution. Here, we extend support graph theory to block-structured matrices and derive eigenvalue bounds for the preconditioned linear system. This extension is crucial for applying support graph theory to the matrices arising in agent-based models. In Section 5, we present our numerical results. We demonstrate the effectiveness of our preconditioner using a series of numerical benchmarks. These experiments validate our theoretical findings and showcase the practical benefits of our approach. Finally, Section 6 concludes the paper with a discussion of our findings, and its implications for real-world use cases. We situate our contributions in the broader theoretical landscape and discuss connections to the existing literature on this problem. Additionally, we provide potential directions for future research in this area.
2. Preliminaries
@@ 1. Step 1:
    Broad-phase collision detection. Identify possible cell contact pairs $(i,j)$. Efficient divide-and-conquer algorithms, such as axis-aligned bounding boxes, are commonly used [39].
@@ 3. Step 3:
    Assemble the friction matrices. Construct the $3\times 3$ cell-cell and cell-substrate friction matrices. The cell-cell friction matrix between cells $i$ and $j$ is given by
    (1) $\Gamma_{ij}^{cc}=A_{ij}\left(\gamma_{\parallel}\mathbf{u}_{ij}\otimes\mathbf{u}_{ij}+\gamma_{\perp}\left(I-\mathbf{u}_{ij}\otimes\mathbf{u}_{ij}\right)\right),$
    where $\gamma_{\parallel}$ and $\gamma_{\perp}$ are the parallel and perpendicular coefficients of friction, respectively (both of which are positive), and $\mathbf{u}_{ij}\in\mathbb{R}^{3}$ is the unit contact vector [26]. If $\mathbf{r}_{i}$ is the position of cell $i$, then
    $\mathbf{u}_{ij}=\frac{\mathbf{r}_{j}-\mathbf{r}_{i}}{\lVert\mathbf{r}_{j}-\mathbf{r}_{i}\rVert}.$
@@ 4. Step 4:
    (2) $\Gamma^{cs}_{i}\mathbf{v}_{i}+\sum_{A_{ij}>0}\Gamma^{cc}_{ij}(\mathbf{v}_{i}-\mathbf{v}_{j})=\mathbf{F}_{i},$
    where $\mathbf{v}_{i}$ is the velocity vector and $\mathbf{F}_{i}$ is the total nonfrictional force acting on the cell. We interpret Equation (2) as one row of a large linear system $\Gamma\mathbf{v}=\mathbf{F}$. For $n$ cells, the friction matrix $\Gamma$ is $3n\times 3n$, and we show that $\Gamma$ is symmetric positive definite. Hence, the conjugate gradient method is an efficient choice for obtaining an accurate solution.
@@ 5. Step 5:
    where the step-size $\Delta t$ is chosen according to the Euler’s method stability criterion. Higher-order integration methods are rarely used. Two exceptions are PhysiCell [14], which employs a second-order Adam’s-Bashford method with a fixed time-step, and [5], where an embedded Runge-Kutta-23 method with an adaptive time-step is employed.
Since the friction matrix $\Gamma$ is composed of $3\times 3$ blocks, we establish a few simple properties of these smaller friction matrices. Note that each cell-cell friction matrix (and the cell-substrate friction matrix of an ellipsoidal cell) is the sum of two orthogonal projectors. Let $\gamma_{\mathrm{max}}$ and $\gamma_{\mathrm{min}}$ be the maximum and minimum elements of $\{\gamma_{\parallel},\gamma_{\perp}\}$, respectively.
Lemma 2.1.
@@
  4. (4)
    its operator norm (with respect to the 2-norm) is $\lVert\Upsilon\rVert=\gamma_{\mathrm{max}}$; and
  5. (5)
@@
Proof.
The projection matrices $\mathbf{u}\otimes\mathbf{u}$ and $I-\mathbf{u}\otimes\mathbf{u}$ are clearly symmetric. Note that $(\mathbf{u}\otimes\mathbf{u})\mathbf{u}=\mathbf{u}$, so $\Upsilon\mathbf{u}=\gamma_{\parallel}\mathbf{u}$. Since the eigenvectors of a real symmetric matrix are orthogonal, take $\mathbf{w}$ such that $\mathbf{w}^{T}\mathbf{u}=0$. Then $(\mathbf{u}\otimes\mathbf{u})\mathbf{w}=0$, so $\Upsilon\mathbf{w}=\gamma_{\perp}\mathbf{w}$. The vector $\mathbf{w}$ lies in the orthogonal complement of $\mathbf{u}$, which is two-dimensional. Both of the eigenvalues are positive, which gives positive definiteness. The operator norm of a symmetric positive definite matrix is the maximum eigenvalue. ∎
Ill-conditioned cell-cell friction matrices will make our preconditioners less effective even if $\Gamma$ as a whole is well-conditioned. In the extreme case of rank-deficient friction matrices, perhaps representing freely rotating objects, our condition number bounds in Section 4 do not hold and support graph preconditioners are likely a poor choice.
@@
Definition 2.2 (Matrix-weighted graph).
A matrix-weighted graph is a triple $G=(V,E,w)$ where $(V,E)$ is an undirected graph with a matrix-valued weight function $w\colon V\times V\to\mathbb{R}^{d\times d}$. We require that $w(e)$ is positive definite for all $e\in E$, that $w(v,v)$ is positive semidefinite for all $v\in V$, and that $w(u,v)=0$ otherwise.
@@
Definition 2.3 (Block Laplacian).
A block Laplacian is a symmetric block matrix whose off-diagonal blocks are either zero or negative definite, and whose block row and column sums are positive semidefinite. For a matrix-weighted graph $G=(V,E,w)$, the block Laplacian of $G$ is the matrix $L$ where
@@
Note that block Laplacians are not diagonally dominant in general, but the condition we impose on their block row and column sums is analogous. We show in Section 4 that this generalized notion of block diagonal dominance is sufficient to apply support graph preconditioners.
Another mathematical structure that seems closely related to the block Laplacian at first glance is the connection Laplacian introduced by Singer and Wu in [37]. However, a key difference between the two structures is that connection Laplacians with real entries require the edge weights to be orthogonal, whereas block Laplacians only require positive definiteness. It is almost never the case that the collision graph yields a connection Laplacian because this would require both friction coefficients to be 1. The solver presented in [24] works on a class matrices satisfying a certain definition of block diagonal dominance which connection Laplacians belong to but block Laplacians generally do not. So the theory of connection Laplacians is not generally applicable to block Laplacians.
@@
Proof.
Let $A$ be a block Laplacian. For all $\mathbf{x}$,
@@
This implies that each term in the summation above is nonnegative, which yields the desired result. ∎
3. Support Graph Preconditioners